# Simpliflow: A Lightweight Open-Source Framework for Rapid Creation and Deployment of Generative Agentic AI Workflows

### v1

@@ I Introduction
Existing agentic AI toolkits—such as LangChain [1], AutoGen [2], and BabyAGI [3]—unlock powerful capabilities (multi-agent collaboration, tool use, memory), but they often demand substantial setup and expertise and tend to execute in non-deterministic ways that complicate debugging, reproducibility, and onboarding for fast experimentation.
@@
The primary contribution of this paper is the introduction of simpliflow [4]. simpliflow is an Open-Source Modular Framework for Generative AI Agentic Workflow Orchestration. Distributed as a Python package (pip install simpliflow) it supports 100+ LLM models/vendors through LiteLLM [5] and allows developers to spin up chains of LLM-powered agents interwoven with custom user-defined functions in seconds. Simpliflow dramatically reduces the time, effort, and specialized knowledge required to incorporate generative agentic AI into applications and workflows at scale. It runs on the IDE of your choice (VSCode, Pycharm, Spyder, Jupyter Notebooks, etc.) and on your platform/OS (Windows, Linux, Mac). By defining a workflow in a simple JSON configuration, users can orchestrate an infinitely long sequence of agent steps with minimal coding. This lowers the barrier to adoption for AI automation: even complex multi-step tasks can be automated or accelerated without building a bespoke agent system from scratch. It presents a lightweight alternative that trades some autonomy for clarity and control: it models workflows as deterministic, linear FSMs with a single, predictable transition per step, avoiding the confusing chat/agent semantics and API fragility reported in other frameworks. simpliflow’s JSON-based workflow definition makes orchestration transparent and auditable, while the engine enforces stepwise sequencing, supports human-in-the-loop approvals, and logs every interaction as structured JSON for inspection. The framework’s nature effortlessly allows AI by AI for AI paradigm i.e. design-time AI support to create AI that can do AI. AI (agents) by AI (in-IDE AI) to do AI (final task) can help you write and execute intelligent machine learning and AI programs using just a high-level intent in natural language i.e. English - all while controlling the creativity, factual nature and diversity of the outputs. It embraces pluggable postprocessor functions—from simple formatters and validators to “AI-to-Action” code execution—so developers can inject domain logic or trigger real actions without leaving the flow. This AI-to-Action feature (code execution) can actually execute the results of the agentic interactions in-situ to for e.g. convert your laptop to a music synthesizer and player or even create and execute quantum programs on a real remote Quantum Computer – all from within the framework.
@@
In the following, we detail Simpliflow’s system architecture and key features, usage, examples and compare it to other prominent agentic AI frameworks (LangChain [1], AutoGen [2], BabyAGI [3], CrewAI [6], SuperAGI [7], etc.)
@@ II System Architecture
  • Agent Class: Simpliflow provides an Agent class (in smartagents.py) which serves as a blueprint for creating agents. Each agent represents a single step in the workflow. An agent is defined by attributes such as its name, role, and task description (prompt), and it encapsulates the logic to communicate with an LLM. Internally, the Agent class leverages LiteLLM [5] – a lightweight library – to interface with various LLMs (OpenAI GPT series, Anthropic Claude, Google models, etc.) behind a common API. This design abstracts away the specifics of different model providers, allowing multi-vendor and multi-model integration transparently. When executed, an agent invokes an LLM with its prompt, producing an output.
@@
Github Repository simpliflow-usage [8], comes with many readymade workflows from various domains. Users can run these or customize these or create their own workflows. A sample of these workflows is shown in Fig. 2.
Figure 2: Example workflows provided in the ”Workflows” directory of the simpliflow-usage repository [8].
The workflows can be crafted by hand or using assistance from In-IDE AI’s like in GitHub Copilot, Cursor etc. or using the Interactive Workflow JSON Generator utility (Fig. 3) that comes along with simpliflow in the simpliflow-usage repo [8].
@@ III Usage, Execution, Accessing Results and Visualization
After the flow execution finishes, the individual agent outputs i.e. intermediate outputs as well as the final result of the flow can be obtained as outputs of the function call to execute the flow. They, as well as the agents, preprocessor function outputs are logged as interactions under the ’Interactions’ directory as JSONs and hence can be visualized easily. Here we visualize them using an In-IDE visualizer JSON Crack [9]. Fig. 9, Fig. 10, and Fig. 11 show visualizations of different workflow structures. Fig. 12 and Fig. 13 show the corresponding logged interactions for two of these workflows.
@@ IV-B Support
  • Supports 100+ LLM Vendors and Models through LiteLLM [5] like Anthropic Claude, Google Gemini, OpenAI ChatGPT, Deepseek, HuggingFace etc.
@@ IV-C Debugging
  • Visualize both agents and interactions using optional in-IDE JSON plugins like JSON Crack [9] or other JSON formatters
@@ V Discussion and Comparison with Existing Frameworks
Several frameworks and toolkits have emerged to help developers build AI agent systems. Here we compare simpliflow with several prominent ones – LangChain [1], AutoGen [2], BabyAGI [3], CrewAI [6], and SuperAGI [7]—focusing on architectural and functional differences. A detailed comparison is provided in Table Simpliflow: A Lightweight Open-Source Framework for Rapid Creation and Deployment of Generative Agentic AI Workflows.
@@ V-A Extremely Easy Setup and Minimal Boilerplate
Frameworks such as AutoGen [2], LangChain [1], and BabyAGI [3] are powerful but typically require more setup time and expertise. Simpliflow addresses this by trading some complexity for simplicity and rapid prototyping. In several agent frameworks (e.g., AutoGen), multiple chat semantics can be confusing. Simpliflow aims to remain simple and abstracts much of the complexity away from developers with a single template.
@@ VI Use Cases/Applications
As explained earlier, simpliflow can help you create and deploy AI agents that can perform diverse tasks across domains. It comes with a collection of example workflows demonstrating its applicability across diverse domains. Below are a few examples included with simpliflow (all available in the simpliflow-usage repository [8]):
@@ VII Future Work and Conclusion
simpliflow presents an alternative in the landscape of agentic AI frameworks. By adopting a declarative JSON-based approach to define linear deterministic workflows, it significantly reduces the barrier to entry for developers and enables unprecedented speed in prototyping and deployment. Its modular architecture, model agnosticism via LiteLLM [5], and powerful ”AI-to-action” capabilities make it a versatile tool for automating a wide array of sequential tasks. Simpliflow is an active project, and several enhancements are planned to expand its capabilities while retaining its core simplicity. Future work will focus on features that improve the robustness, performance, and usability of the framework for, e.g., Enhanced Error Handling, and Async and Parallel Agent Execution.

### v2

@@ I Introduction
Existing agentic AI toolkits—such as LangChain [langchain_github], AutoGen [autogen_github], and BabyAGI [babyagi_github]—unlock powerful capabilities (multi-agent collaboration, tool use, memory), but they often demand substantial setup and expertise and tend to execute in non-deterministic ways that complicate debugging, reproducibility, and onboarding for fast experimentation.
@@
The primary contribution of this paper is the introduction of simpliflow [simpliflow_github]. simpliflow is an Open-Source Modular Framework for Generative AI Agentic Workflow Orchestration. Distributed as a Python package (pip install simpliflow) it supports 100+ LLM models/vendors through LiteLLM [litellm_github] and allows developers to spin up chains of LLM-powered agents interwoven with custom user-defined functions in seconds. Simpliflow dramatically reduces the time, effort, and specialized knowledge required to incorporate generative agentic AI into applications and workflows at scale. It runs on the IDE of your choice (VSCode, Pycharm, Spyder, Jupyter Notebooks, etc.) and on your platform/OS (Windows, Linux, Mac). By defining a workflow in a simple JSON configuration, users can orchestrate an infinitely long sequence of agent steps with minimal coding. This lowers the barrier to adoption for AI automation: even complex multi-step tasks can be automated or accelerated without building a bespoke agent system from scratch. It presents a lightweight alternative that trades some autonomy for clarity and control: it models workflows as deterministic, linear FSMs with a single, predictable transition per step, avoiding the confusing chat/agent semantics and API fragility reported in other frameworks. simpliflow’s JSON-based workflow definition makes orchestration transparent and auditable, while the engine enforces stepwise sequencing, supports human-in-the-loop approvals, and logs every interaction as structured JSON for inspection. The framework’s nature effortlessly allows AI by AI for AI paradigm i.e. design-time AI support to create AI that can do AI. AI (agents) by AI (in-IDE AI) to do AI (final task) can help you write and execute intelligent machine learning and AI programs using just a high-level intent in natural language i.e. English - all while controlling the creativity, factual nature and diversity of the outputs. It embraces pluggable postprocessor functions—from simple formatters and validators to “AI-to-Action” code execution—so developers can inject domain logic or trigger real actions without leaving the flow. This AI-to-Action feature (code execution) can actually execute the results of the agentic interactions in-situ to for e.g. convert your laptop to a music synthesizer and player or even create and execute quantum programs on a real remote Quantum Computer – all from within the framework.
@@
In the following, we detail Simpliflow’s system architecture and key features, usage, examples and compare it to other prominent agentic AI frameworks (LangChain [langchain_github], AutoGen [autogen_github], BabyAGI [babyagi_github], CrewAI [crewai_github], SuperAGI [superagi_github], etc.)
@@ II System Architecture
  • Agent Class: Simpliflow provides an Agent class (in smartagents.py) which serves as a blueprint for creating agents. Each agent represents a single step in the workflow. An agent is defined by attributes such as its name, role, and task description (prompt), and it encapsulates the logic to communicate with an LLM. Internally, the Agent class leverages LiteLLM [litellm_github] – a lightweight library – to interface with various LLMs (OpenAI GPT series, Anthropic Claude, Google models, etc.) behind a common API. This design abstracts away the specifics of different model providers, allowing multi-vendor and multi-model integration transparently. When executed, an agent invokes an LLM with its prompt, producing an output.
@@
Github Repository simpliflow-usage [simpliflow_usage_github], comes with many readymade workflows from various domains. Users can run these or customize these or create their own workflows. A sample of these workflows is shown in Fig. 2.
Figure 2: Example workflows provided in the ”Workflows” directory of the simpliflow-usage repository [simpliflow_usage_github].
The workflows can be crafted by hand or using assistance from In-IDE AI’s like in GitHub Copilot, Cursor etc. or using the Interactive Workflow JSON Generator utility (Fig. 3) that comes along with simpliflow in the simpliflow-usage repo [simpliflow_usage_github].
@@ III Usage, Execution, Accessing Results and Visualization
After the flow execution finishes, the individual agent outputs i.e. intermediate outputs as well as the final result of the flow can be obtained as outputs of the function call to execute the flow. They, as well as the agents, preprocessor function outputs are logged as interactions under the ’Interactions’ directory as JSONs and hence can be visualized easily. Here we visualize them using an In-IDE visualizer JSON Crack [jsoncrack_github]. Fig. 9, Fig. 10, and Fig. 11 show visualizations of different workflow structures. Fig. 12 and Fig. 13 show the corresponding logged interactions for two of these workflows.
@@ IV-B Support
  • Supports 100+ LLM Vendors and Models through LiteLLM [litellm_github] like Anthropic Claude, Google Gemini, OpenAI ChatGPT, Deepseek, HuggingFace etc.
@@ IV-C Debugging
  • Visualize both agents and interactions using optional in-IDE JSON plugins like JSON Crack [jsoncrack_github] or other JSON formatters
@@ V Discussion and Comparison with Existing Frameworks
Several frameworks and toolkits have emerged to help developers build AI agent systems. Here we compare simpliflow with several prominent ones – LangChain [langchain_github], AutoGen [autogen_github], BabyAGI [babyagi_github], CrewAI [crewai_github], and SuperAGI [superagi_github]—focusing on architectural and functional differences. A detailed comparison is provided in Table VII.
@@ V-A Extremely Easy Setup and Minimal Boilerplate
Frameworks such as AutoGen [autogen_github], LangChain [langchain_github], and BabyAGI [babyagi_github] are powerful but typically require more setup time and expertise. Simpliflow addresses this by trading some complexity for simplicity and rapid prototyping. In several agent frameworks (e.g., AutoGen), multiple chat semantics can be confusing. Simpliflow aims to remain simple and abstracts much of the complexity away from developers with a single template.
@@ VI Use Cases/Applications
As explained earlier, simpliflow can help you create and deploy AI agents that can perform diverse tasks across domains. It comes with a collection of example workflows demonstrating its applicability across diverse domains. Below are a few examples included with simpliflow (all available in the simpliflow-usage repository [simpliflow_usage_github]):
@@ VII Future Work and Conclusion
simpliflow presents an alternative in the landscape of agentic AI frameworks. By adopting a declarative JSON-based approach to define linear deterministic workflows, it significantly reduces the barrier to entry for developers and enables unprecedented speed in prototyping and deployment. Its modular architecture, model agnosticism via LiteLLM [litellm_github], and powerful ”AI-to-action” capabilities make it a versatile tool for automating a wide array of sequential tasks. Simpliflow is an active project, and several enhancements are planned to expand its capabilities while retaining its core simplicity. Future work will focus on features that improve the robustness, performance, and usability of the framework for, e.g., Enhanced Error Handling, and Async and Parallel Agent Execution.
TABLE I: Comparison of AI Orchestration Frameworks
Feature / Aspect Simpliflow (proposed) LangChain AutoGen BabyAGI CrewAI SuperAGI
Orchestration Style Predefined linear workflow (FSM-like sequence) – fully deterministic flow control by design. Flexible chains and dynamic agents (LLM-driven decisions); mix of static pipelines and agent loops. Dynamic multi-agent chat framework; no built-in fixed process (flows emerge from agent interactions). Autonomous loop generates and executes tasks continuously until stopped (emergent sequence, not predetermined). Supports both explicit Flows (event-driven steps) and agent Crews (autonomous team), allowing structured or dynamic behavior. Typically dynamic agent planning with toolkit; user defines goal, agent decides steps (though can script sequences via GUI config).
Determinism Yes: Execution path is fixed by JSON (only LLM output content varies). Promotes reproducibility and predictability. Partial: Deterministic if using simple Chains; Not if using Agent mode (outcome path can vary run to run). No: Agents decide next moves in conversation; flow can diverge. Hard to reproduce exact trajectories without forcing a script. No: Inherently non-deterministic (self-generated tasks differ each run, depending on outcomes and memory). Yes/Partial: Flows can be deterministic; Crews introduce non-determinism via agent autonomy. Tends toward determinism when using flow control explicitly. No: Agents often use AI reasoning to choose actions; results may differ each run. Some determinism if using fixed sequence mode in UI, but core design assumes autonomy.
Primary Abstraction Workflow JSON with agents & transitions; minimal code (just run it via simpliflow API). Library/API with classes (Chains, Agents, Tools, Memory, etc.); requires writing Python/JS code or YAML configurations. Library/API with agent classes; define agents and launch chats programmatically (some no-code UI add-on available). Script/Template – typically implemented as Python script demonstration; not a full library interface (often integrated into others). Library + Platform: Code library for Flows/Crews and optional control plane UI for enterprise; requires coding to define roles and flows, plus optional configuration in UI. Platform (GUI + YAML) – offers a web interface and config files to set up agents, with underlying Python framework. Focus on less coding, more configuration.
Learning Curve Low: Very simple syntax and concept (JSON steps). Quick to get running with provided examples. Overall developer-friendly. Moderate–High: Steep for newcomers – many concepts (prompts, memory, agents, etc.) and rapidly evolving APIs. Good documentation and community, but complexity is inherent to flexibility. Moderate: Need understanding of async multi-agent paradigms. The concept of agents chatting is intuitive, but mastering advanced features (teaching agents, custom tools) adds complexity. Low (for basic use): The original code is short and easy to run. High to modify or extend – since not designed as extensible framework, customizing behavior or adding tools requires significant coding. Moderate: Core ideas (agents = crew members, tasks = flows events) require some learning. The framework is lean but one must write code to utilize it fully. Enterprise features (like deployment, telemetry) add to learning but are optional. Low–Moderate: The GUI and templates make basic agent setup easy (non-coders can use it). However, fully leveraging it (custom tools, debugging agent decisions) requires understanding the underlying logic.
Modularity/Extensibility High: Very modular – plug in any LLM via LiteLLM (100+ models), add custom postprocessor functions (for new actions or integrations) easily. Designed to fit into any project (supports many tools, data, etc.). High: Provides many extension points (custom tools, custom memory, new chain types). Has a large ecosystem of integrations (vector DBs, model APIs). Coupled somewhat to its abstractions (one must conform to LangChain interfaces when extending). High: Allows custom agents, custom tool functions, and integration of human agents. Primarily extendable through writing Python subclasses or functions. Multi-language support (it has .NET version) extends reach. Low: Not modular in design – it’s a specific loop structure. Extensions usually involve merging with another framework (e.g., adding LangChain tools) rather than BabyAGI itself providing extension hooks. High: Built for customization – users can craft unique crew compositions, define low-level agent behaviors, and integrate any model (including local). Also offers plugin tool set (crewai[tools]). Somewhat complex to extend due to need for code, but very flexible in capable hands. High: Many built-in tool integrations and memory. Allows adding new “skills” or tools via its plugin system. You can swap in different vector DBs, models, etc. Extensible, though one might need to follow the framework’s conventions.
Memory & State Management Minimalist: No built-in long-term memory between agents beyond passing the immediate output to next. Each run is fresh (unless user manually feeds prior logs in). Relies on external means if persistent memory needed (BYO vector DB for RAG if required). State is primarily the JSON structure and the in-memory variables during execution. Rich: Provides memory classes for chat history, knowledge graph, etc. Easily add context memory or tracking of state across steps. Also supports caching of LLM calls to reuse results. State management is a key part of LangChain’s value prop (but adds complexity). Conversation history is inherently the state (messages exchanged). Also supports function calling where intermediate results can be stored. No global memory beyond the chat unless user adds (though one could integrate external memory). Some focus on LLM self-evaluation to update state (teaching agents). Central: Uses a vector store as memory to remember completed tasks and results, enabling the loop to “remember” what’s done. The task list is state in memory. No concept of ephemeral per-step state except through that persistent memory. Offers both transient memory (agent context that persists during a crew’s run) and options for persistent memory (delegation to external stores). Telemetry can log state of each agent. State management is explicit: the developer can choose what each agent sees/retains. CrewAI can run agents sequentially or in parallel threads, so synchronization of state is up to flow design. Yes: Typically integrates with a vector database or other memory to allow agents to store and retrieve information over long tasks. State can include past tool outputs, conversation, etc., managed by the framework. The platform keeps logs and agent state for monitoring.
Human Oversight Yes: Built-in via require_human_approval. Human can intercept at specific steps. Otherwise, user is not involved during run (one-shot execution unless paused). Indirect: No specific “pause for approval” feature; but developers can insert manual steps or use an AgentTool that represents a human. LangChain mostly assumes either fully automated or externally supervised outside the chain. Yes: AutoGen can include a human as one of the agents (e.g., an interactive console agent) to provide oversight or input in the multi-agent chat. Not a one-click setting, but supported conceptually (human-in-the-loop as an agent). No: Not in original design – it’s autonomous until stopped. A user could monitor the console and intervene by aborting or editing code, but no formal mechanism for stepwise approval. Not directly a one-step toggle; however, one can design a Crew with a human “agent” role or require certain steps to output to console for verification. CrewAI is more about full automation, assuming thorough testing rather than live human interjection. Somewhat: Through the GUI, a user might inspect agent decisions in real-time and could adjust configurations. But once an agent is running, it’s generally autonomous. No explicit pause/approve feature standard, though enterprise usage may involve a human supervisor restarting or adjusting if the agent stalls.
Notable Strengths Simplicity & Determinism: Very easy to use and integrate; predictable linear logic; quick to adapt to many domains with little overhead. Great for well-defined process automation with LLMs. Extensiveness & Community: Huge array of features and integrations; standard choice for complex LLM apps. Can handle complex dialogues, tool chains, and has active support and improvements. Multi-agent Collaboration: Enables sophisticated interactions between AI agents (and humans) – good for scenarios where agents brainstorm or critique each other. Offers advanced research features (teaching, self-tuning) and supports cutting-edge use cases in agent research. Autonomy & Creativity: Maximizes AI autonomy; useful for exploring open-ended problem solving. Minimal setup to witness an agent that self-generates tasks. Inspired many derivatives (e.g., integrated into other frameworks). Speed & Control: High-performance execution, independent (no LangChain bloat). Combines structured flows with autonomous agent teams – versatile for simple or complex tasks. Enterprise readiness (monitoring, security) built-in, making it suitable for production pipelines where both AI freedom and oversight are needed. Feature-Complete & User-Friendly: Provides a lot out-of-the-box (UI, memory, tools, safety checks). Lower barrier for non-developers to deploy an AI agent solution. Designed for scalability and robustness in production (logging, fail-safes, etc.).
Notable Trade-offs Limited Autonomy: Not suited for scenarios where the AI needs to figure out the process itself or react outside a fixed script. Lacks built-in long-term memory and learning – relies on static design and user iteration to improve. Complex & Heavy: Using LangChain for simple tasks can be overkill. The system can become complex with many moving parts; debugging a large LangChain agent can be difficult. Performance can suffer if not carefully managed (overhead of many components). Orchestration Overhead: Without a fixed process control, managing a large multi-agent workflow might require custom logic. Potentially unpredictable outcomes require careful testing. Also, being relatively new, it’s evolving (Microsoft has even announced a next-gen Agent Framework), so long-term support is something to watch. Lacks Structure/Safety: Can easily go off-track or produce nonsensical tasks if the LLM errs. Not directly enterprise-ready; needs to be combined with other tools for things like monitoring or tool usage. Essentially a proof-of-concept made practical only when embedded in a larger system. Requires Coding & Setup: Despite its power, one still needs to write and orchestrate code to define crews/flows. The enterprise features (AMP) might need infrastructure. If only a simple linear chain is needed, CrewAI could be unnecessarily complex relative to simpliflow. Also, being newer, its community is smaller than LangChain’s (though growing). Complex Under the Hood: The simplicity for the user is achieved by a complex underlying system. Debugging why a SuperAGI agent did something might be non-trivial. Also, running a persistent agent server might be resource-intensive. There’s some lock-in to how SuperAGI structures problems (less developer control over each decision).